Assume a borrower wants to borrow $275,000 for 30 years at 4.50 %. first two monthly payments for a CAM and for a CPM? (4) 14 answer all 3 please
Answer #1

14. CPM ( Constant Payment Mortgage)

In this constant payment is calculated over the period of the loan

Total Loan = $275000

Time Period = 30 years or 30*12 = 360 months

Interest Rate = 4.50% annually or 4.50/12 = 0.375 % monthly

Formula

Let monthly payment be X

275000 = [(X/0.00375)*(1-(1/(1.00375)360))]

X = (275000*0.00375)/(1-(1/(1.00375)360))

X = 1031.25/0.7401 = $1393.38

CAM ( Constant Amortization Mortgage )

In this Principal each month remain constant and interest for the period is added

Principal = 275000/360 = $763.88 per month

For First Month

Interest = 275000*0.00375 = $1031.26

Payment for first month = 763.88 +1031.26 = $1795.14

For Second Month

Interest = (275000-763.88)*0.00375 = 274236.12*0.00375= $1028.39

Payment for second month = 763.88 +1028.39 = $1792.27

hence answer is d